The commercial lead acid battery has 15 plates the number of negative plates will be
1)9
2)8
3)7
4)10
1
Expert's answer
2019-12-16T11:10:07-0500

The number of negative plates is more than positive plates in lead acid battery

Number of negative plates

= (n+1)/2

= (15+1)/2

= 8

No of positive plates

= (n-1)/2

= (15-1)/2

= 7

hence the number of negative plates will be 8